OpenStudy (anonymous):

Write the equation 2x - 3y = 6 in slope-intercept form. y = x + 3 y = x + 3 y = x - 2 pls help

OpenStudy (anonymous):

You got to get the y by its self on one side and then solve.

OpenStudy (beautifulmystery.):

To write in slope-intercept form means to solve for the variable y. 2x - 3y = 6 Subtract 2x from each side -3y = -2x + 6 Divide everything by -3 -3y/-3 = -2x/-3 + 6/-3 Simplify y = (2/3)x - 2

OpenStudy (beautifulmystery.):

y = x - 2

OpenStudy (anonymous):

thanks